OpenStudy (anonymous):

The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (anonymous):

heyy

random231 (random231):

okay first of all parallel lines have the same slope. so slope of the required line is -2. now use: y-5=(-2)(x-4)

OpenStudy (anonymous):

is that what i use to plug in

random231 (random231):

so arent u getting the answer?

OpenStudy (anonymous):

because im not sure how to solve the equation and the answer choices are y = −2x + 13 y = negative 1 over 2x + 7 y = 1 over 2x + 3 y = − 2x − 3

random231 (random231):

its the first one

OpenStudy (anonymous):

thank you so much i appreciate it